OpenStudy (anonymous):

I understand how to solve now. But how would I used a calculator to do all of that? haha

OpenStudy (agent47):

f'(x)=3x^2+4x-8 Set that equal to 0 and solve for x: 3x^2+4x-8=0, x=-2.431, or x=1.097

OpenStudy (anonymous):

oh okay [:

OpenStudy (agent47):

b) Find the 2nd derivative: 6x+4 Plug in x=-2.431: 6*(-2.431)+4<0, so x=-2.431 is a max

OpenStudy (agent47):

Plug in x=1.097 6*1.097+4>0, so x=1.097 is a min

OpenStudy (blockcolder):

I think the function in the problem is f'(x), not f(x).

OpenStudy (anonymous):

Thanks a bunch, I'll review this.